OpenStudy (anonymous):

what is the first 3 terms of the geometric sequence and the nth term of: a1=4, r=1/2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

4 , 2, 1 nth term = a r ^ (n-1) = 4 * (1/2^(n-1)

OpenStudy (amistre64):

a1, 4 a2, 4(1/2) a3, 4(1/2)(1/2) a4, 4(1/2)(1/2)(1/2)

OpenStudy (amistre64):

an = 4(1/2)^(n-1) ; yep

OpenStudy (anonymous):

\[4(1/2)^n-1\]

OpenStudy (amistre64):

\[a_n=4(1/2)^{n-1}\]

OpenStudy (amistre64):

gotta put the {} around the exponent in latex to get it all